A father is three times as old as his son. After fifteen years the father will be twice as old as his son's age at that time. Hence the father's present age is
Found 2 solutions by mananth, karthiktcs:
Answer by mananth(16946)

Let son's age = x
father's age =3x
..
After 15 years
son's age = x+15
father's age = 3x+15
..
3x+15=2(x+15)
3x+15=2x+30
x=15.. son's age now
Father's age now = 3x = 45

Answer by karthiktcs(3)

let the present age of son be x
and the present age of father be 3x
after fifteen years
(3x+15)=2(x+15)
(3x+15)=(2x+30)
3x+15-2x-30=0
x-15=0
x=15
father age will be 3*15=45